OpenStudy (anonymous):

How much is 2/3 cup plus 2/3 cup. I think it is 1 1/3 correct or no? If not correct please tell me what need to do not telling me the answer I want to do it on my own.

OpenStudy (superflytyguy):

You are right. 2/3 + 2/3 = 4/3 which is an improper fraction. 3/3 = 1 4/3 = 1 1/3

OpenStudy (anonymous):

Thanks! Closing down :( Ihavefinishmypracticequestion

OpenStudy (uri):

Yeah you're correct.